OpenStudy (anonymous):

you invest $3,000 in an account with an interest rate of 5.5% compounded continuously. How much money would be in the account after 5 years? Round your answer to the nearest whole number.

myininaya (myininaya):

compounded continuously means you will be using this formula: \[A=A_0e^{rt}\]

OpenStudy (anonymous):

There.

myininaya (myininaya):

A_0 is the initial investment r is the rate t is number of years A is the amount after time t

myininaya (myininaya):

Ok...I was hoping to involved the asker some in the question.

OpenStudy (anonymous):

You both have completely different equations…. now I'm confused.

OpenStudy (anonymous):

I got my answer from Algebra.com :/ I suppose it could be wrong, just have myininaya help you out then you decide which one you wanna go with.

OpenStudy (anonymous):

okay… i understand the equation.. but where do i plug in what. @myininaya

myininaya (myininaya):

equations are exactly the same we just have different letters (they have the same meaning though)

myininaya (myininaya):

his F is my A his P is my A_0

OpenStudy (anonymous):

would it be a_0 =3,00 r = 5.5 and t = 5?

OpenStudy (anonymous):

3,000*

myininaya (myininaya):

yes 3000 is the initial investment r=0.055 t=5

myininaya (myininaya):

r=0.055 because we had r was 5.5% not 5.5

myininaya (myininaya):

percent means /100

myininaya (myininaya):

5.5/100=0.055=r

OpenStudy (anonymous):

ohh. okay, i understand that.. so then id take a=3,000^(0.055)(5) ?

myininaya (myininaya):

close you are missing one little thing

myininaya (myininaya):

\[A=3000e^{0.055 \cdot 5}\]

OpenStudy (anonymous):

whats the e for…?

myininaya (myininaya):

\[A=A_0e^{rt}\] e is the base you use if is compounded continuously

myininaya (myininaya):

if it isn't compounded periodically we would use \[A=A_0(1+\frac{r}{n})^{nt}\] where n is the number of compounding periods per year

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}A_0(1+\frac{r}{n})^{n t}=A_0e^{rt}\]

myininaya (myininaya):

you don't need to know that last thing

myininaya (myininaya):

but it why we use e

myininaya (myininaya):

the explanation is a calculus topic

myininaya (myininaya):

and I made a type-o above if it compounded periodically* (not isn't )

myininaya (myininaya):

if it is compounded periodically we would use \[A=A_0(1+\frac{r}{n})^{nt}\] where n is the number of compounding periods per year *